OpenStudy (anonymous):

when i do an exponential function equation thats based on time, how would i change the time from months to weeks?

OpenStudy (jdoe0001):

you'd would use the unit of conversion, usually the figure will be expressed as some fraction like say mi/hr or angles/sec or gallons/mi so you'd use the fraction correspondent to the unit you want changed, and put the unit you want cancelled in the opposite spot say you have 3months to weeks, well we know that there are 4 weeks in 1 month so to cancel out months, you'd put months as denominator, thus \(\bf 3month\times \cfrac{4week}{month}\implies \cfrac{3\cancel{month}\times 4week}{\cancel{month}}\) if instead, you wanted weeks to months, then you'd flip the same unit correspondent elements, say 3weeks to months, \(\bf 3week\times \cfrac{month}{4week}\implies \cfrac{ 3\cancel{week}\times month}{4\cancel{week}}\)